OpenStudy (anonymous):

Help with Vectors question? Consider the following lines: L1: [x, y] = [3, -2] + t[4, -5] L2: [x, y] = [1 , 1] + s[7 , k] a) For what value of k are the lines parallel? b) For what value of k are the lines perpendicular? Please show all work. Thanks :)

OpenStudy (anonymous):

Hey Callisto :)

OpenStudy (callisto):

Hmm.... what is t and s?!

OpenStudy (anonymous):

I dont know :P

OpenStudy (callisto):

That doesn't look like something I've learnt before :( What I've learn is that For two vectors \[m_1i + n_1 j \ and \ m_2i+n_2j\] (i) if they are // \[\frac{m_1}{m_2} = \frac{n_1}{n_2}\] (ii) if they are perpendicular a.b =0 ie \(m_1m_2+n_1n_2=0\) That's all I know :(

OpenStudy (anonymous):

hmm... maybe they are variables?

OpenStudy (amistre64):

s and t just mark the vectors, they scale the vectors to raech all the points on the line

OpenStudy (amistre64):

stretch the first to equal the second and then you will be able to solve for k

OpenStudy (anonymous):

Can you take me through this, Im new to this concept :P

OpenStudy (amistre64):

the t resembles distribution; just push it thru the grouping symbols

OpenStudy (amistre64):

[4t , -5t] = [7 , k] solve for t with the first component and youll know what k has to be

OpenStudy (anonymous):

So, 4t = 7?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

t=7/4

OpenStudy (anonymous):

t=1.75

OpenStudy (amistre64):

7/4 is fine; you can decimate it afterwards if it makes you feel better

OpenStudy (anonymous):

1.75x-5=k?

OpenStudy (anonymous):

k=-8.75

OpenStudy (amistre64):

-5t = k

OpenStudy (amistre64):

i read your times as an x lol

OpenStudy (anonymous):

but t is 7/4 right so -5 times 1.75 right?

OpenStudy (anonymous):

Oh sorry about that :P

OpenStudy (amistre64):

yes, -5 times 1.75 = k for parallel

OpenStudy (anonymous):

How would we do perpendicular?

OpenStudy (amistre64):

callisto had that one called multiply the vectors componentwise; then add the results and equate them to 0

OpenStudy (amistre64):

4,-5 7, k ----- 28 -5k = 0

OpenStudy (amistre64):

its called a dot product if you wanna get technical :)

OpenStudy (anonymous):

So in hat last line we solve for k, and then are we done with the question?

OpenStudy (anonymous):

**that

OpenStudy (amistre64):

yes, unless you have something hiding underneath it :)

OpenStudy (anonymous):

LOL, nope, I dont :P

OpenStudy (anonymous):

THANKS FOR YOUR HELP amistre64! You are a life saver!

OpenStudy (amistre64):

callisto did well too

OpenStudy (callisto):

I did nothing..... I'm still trying to understand the t and s :S

OpenStudy (anonymous):

LOL

OpenStudy (amistre64):

think of the t or s, they had to use 2 different scalars just becasue ... think of them as how far you would stretch a slinky to get it to cover 3 feet or 10 feet or whatever feetage

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

the vector itself gives us a direction to point in; the scalar allow us to hit all the points in that direction to form a line with

OpenStudy (anonymous):

Can you tell me how to write the therefore statements?

OpenStudy (amistre64):

therefore, yada yada ?? im no good with that part im afraid :)

OpenStudy (anonymous):

LOL so just the lines are parallel when k = -8.75 right?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

Ok, THANKS AGAIN + CALLISTO!!

OpenStudy (callisto):

Why did you multiply the number in the bracket by t for t[4, -5], but did nothing for s[7 , k] ?

OpenStudy (anonymous):

hmm.... yea I was wondering about that too, and why is the first part of each line left out?

OpenStudy (anonymous):

@amistre64 Can you please elaborate :)

OpenStudy (amistre64):

we simply were looking to get 4,-5 stretched to the same values as 7,k in order to have an identical vector .... having identical parts. there was no need to alter 7,k in any way

OpenStudy (anonymous):

OHHH, I see noww...

OpenStudy (anonymous):

Thanks :)

OpenStudy (amistre64):

yep

OpenStudy (callisto):

and the things in front of t[.. , ..] and s[.. , ..] are not our concern?

OpenStudy (amistre64):

parallel lines are parallel regardless of what points they are attached to

OpenStudy (amistre64):

the direction of a line in space is determined by its vector part; not its point part. The point just anchors it to a specific spot

OpenStudy (callisto):

Oh... so those were points? :|

OpenStudy (amistre64):

yes; read it as: at this point, anchor this vector, and stretch it to form a line

OpenStudy (callisto):

Wow! Thanks!!!! As I've said, I've never seen such thing before.... Thank you so much!!!

OpenStudy (amistre64):

youre welcome